right triangle.the longest side of a right triangle is 3 inches less than twice the shortest side.the third side measures 12 inches.find the length ot the shortest side
Answer by checkley71(8403)

X^2+12^2=(2X-3)^2
X^2+144=4X^2-12X+9
X^2-4X^2+12X+135=0
-3X^2+8X+135=0
3X^2-8X-135=0 USING THE QUADRATIC EQUATION WE GET:
X=(-B+-SQRT[B^2-4AC])/2A
X=(8+-SQRT[64-4*3*-135])/2*3
X=(8+-SQRT[64+1620])/6
X=(8+-SQRT1684)/6
X=(8+-41.036457)/6
X=(8+41.036457)/6
X=49.036457/6
X=8.17 INCHES FOR THE SHORTEST SIDE.